# A Poisson-Guided Decomposition Network for Extreme Low-Light Image Enhancement

### v1

@@ III-B3 Noise Loss
which is used to suppress the intensity of noise ($\alpha$ is set as $10^{-6}$ computational stability).
Figure 3: Visual comparison with other state-of-the-art low-light enhancement methods: (a) low-light input image, (b) LIME [13], (c) deep Retinex [24], (d) SRIE [45], (e) zero-DCE [46], (f) Proposed, (g) Ground Truth. Notice that our method achieves better quality in the enhanced image than the other methods.
IV Experimental Results
IV-1 Dataset and Experimental Setup
We have used the Low Light dataset (LOL dataset) [24], which contains 500 paired low-light and normal-light images which are collected by changing exposure time and ISO, while other configurations of the cameras are fixed. During training (100 epochs), each input consists of $3\times 128\times 128$ sized RGB color patches of paired low-light and normal-light images.
IV-2 Baseline Methods
To demonstrate the effectiveness of our proposed methods, we compare with the following baseline methods: LIME [13], deep Retinex [24], SRIE [45], Light-weight low-light enhancement [26], and zero-DCE [46].
IV-3 Visual Comparison
We display the visual results of our method along with the baseline ones on two different scenes at varying illumination levels in Figures 2 and 3. In both cases, we have considered input image with four different levels of illumination (reduced as we move down from the first row to the fourth row). Notice that in both Figures 2 and 3, there are color distortions in the output images by the baseline methods from (b) to (f). Our method produces best visual quality restored images without color distortion shown in Fig. 2(f) and Fig. 3(f). Another important fact is that such color distortion gets worse with very low level of illumination (as is the case for last rows in both Figures 2 and 3) which is the case of extreme low-light condition. In such cases, our method significantly outperforms the baseline methods.
IV-4 Quantitative Comparison
We use three referenced and one no-referenced image metrics to quantify the comparative performance evaluation of our method with benchmarks: peak signal-to-noise ratio (PSNR) [48], structural similarity index measure (SSIM) [48], mean absolute error (MAE) [49], and no-reference image quality evaluator (NIQE) [48]. For PSNR and SSIM, a higher value indicates better quality, whereas for NIQE and MAE, a lower value indicates better quality of the enhanced image. In Table I, we summarize the performance of low-light enhancement using our method and the baselines. We found that our method outperforms the existing methods in terms of both PSNR and SSIM [50]. In addition, our method achieves competitive performance with the baseline methods in terms of NIQE and MAE metrics. Experimental results confirm that our method can perform efficient low-light enhancement even in extreme low-light conditions.
Our research demonstrates that the dominance of Poisson noise in extreme low-light images significantly contributed to color distortion, which gets severe with extreme low-lightness. Poisson noise, being signal-dependent, is more prominent in low-light conditions where photon counts ($c$) are low. Visual results confirm our method can perform efficient low-light enhancement even in extremely low-light conditions.
@@ IV-4 Quantitative Comparison
TABLE I: Comparison for different methods. The best are marked as red, the second best are marked as blue.
Method PSNR(dB) SSIM NIQE MAE
LIME [13] 15.66 0.42 9.35 35.57
Deep Retinex [24] 16.06 0.58 4.31 35.78
SRIE [45] 9.85 0.40 7.72 80.00
Zero-DCE [46] 12.35 0.55 3.42 61.71
Light-weight [26] 15.99 0.59 4.54 —-
Proposed 17.12 0.67 3.54 32.33
V Conclusion
@@
In this work, we attempt to address a challenging task of extreme low-light image enhancement and denoising. We demonstrate with experimental finding that many baseline methods suffer from color distortion when the illumination level of the scene/object is extremely poor. We introduced a novel image decomposition framework in contrast to the Retinex model of image decomposition. Further, our encoder-decoder-based deep learning method could produce superior recovery of extreme low-light images without any visible color distortion. Given that Poisson noise dominates in extreme low-light images because of limited photon counts, specialized Poisson denoising techniques are essential for preserving image quality and structural details. In future research, we will focus on developing denoising models that can robustly handle mixed noise distributions while enabling more efficient solutions for real-time applications.

### v2

@@ IV-B3 Noise Loss
which is used to suppress the intensity of noise ($\alpha$ is set as $10^{-6}$ computational stability). Notice that the loss in (15) is inspired by the existence of Poisson noise in extreme low-light images.
We summarize the loss function in Figure 2. Notice that introducing $\mathcal{L}_{n}$ loss is one of the novel contributions in our method. The hyperparameters $\gamma$, $\lambda_{1}$ and $\lambda_{2}$ are set using grid-search method during validation stage. Finally, we refer to our proposed method as Poisson Informed Retinex of Low-light images (PIRL).
@@
TABLE I: Comparison for different methods on LOL dataset. The best are marked as red, the second best are marked as blue.
Method Reference PSNR(dB) SSIM NIQE
LIME [13] IEEE TIP 2016 15.66 0.42 9.35
LIME-filtered [13] IEEE TIP 2016 16.77 0.59 4.53
SRIE [56] IEEE/CVF CVPR 2016 9.85 0.40 7.72
Deep Retinex [24] BMVC 2018 16.06 0.58 4.31
RRM [57] IEEE TIP 2018 13.88 0.658 5.810
HybridNet [58] IEEE TIP 2019 16.60 0.66 3.42
Zero-DCE [44] IEEE/CVF CVPR 2020 12.35 0.55 3.42
STAR [59] IEEE TIP 2020 12.64 0.538 6.205
LR3M [60] IEEE TIP 2020 10.22 0.399 7.522
DRBN [61] IEEE/CVF CVPR 2020 16.75 0.659 4.724
RUAS [62] IEEE/CVF CVPR 2021 16.40 0.583 6.340
RetinexDIP [63] IEEE TCSVT 2021 9.442 0.334 7.070
EFINet [64] IEEE TCSVT 2022 13.456 0.626 - -
SCL-LLE [65] AAAI 2022 12.42 0.520 7.63
SCI [66] IEEE/CVF CVPR 2022 14.78 0.520 7.87
SGZ [67] IEEE/CVF WACV 2022 15.93 0.570 7.82
UEGAN [68] IEEE TIP 2020 12.22 0.262 5.627
DUNP [43] IEEE TCSVT 2022 15.49 0.654 3.73
GDP [42] IEEE/CVF CVPR 2023 15.89 0.542 6.13
NeurBR [41] Pattern Recognition 2024 11.38 0.440 7.50
PnP Retinex [69] IEEE TNNLS 2024 17.12 0.66 3.95
PNAR [37] IEEE SPL 2021 16.53 0.583 - -
IQA [48] IEEE SPL 2022 17.95 0.658 - -
PIRL Proposed 17.12 0.67 3.54
V Experimental Results
V-1 Dataset and Experimental Setup
We used a publicly available Low Light (LOL) image dataset [24] to train and test our deep learning method. The LOL dataset has 500 paired low-light and normal-light images. These images were captured by changing the exposure time and ISO, whereas the other configurations of the cameras are unchanged. The training process employs the model for the joint decomposition of low-light images into illumination, reflectance, and noise components.
During training, each input consists of $3\times 128\times 128$ sized RGB color patches of paired low-light and normal-light images. The ADAM optimizer is utilized with parameter settings: learning rate ($lr$) initialized at $0.001$, stability factor $\epsilon=10^{-8}$, and moving average control coefficients $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The training is conducted for 100 epochs with a fixed batch size of 16. Then, we pre-processed the dataset by cropping the images into patches of $128\times 128$. In addition, the proposed algorithm is implemented on the PyTorch platform.
V-2 Benchmarking Methods
To demonstrate the effectiveness of our proposed method, we compare the results with state-of-the-art methods. We compare with both sub-types of classical and deep learning methods. The non-deep learning (i.e. classical) methods compared are: low-light image enhancement (LIME) [13], LIME-filtered [13], simultaneous reflectance and illumination estimation (SRIE) [56], robust retinex model (RRM) [57], structure and texture aware retinex (STAR) [59], and low-rank regularized retinex model (LR3M) [60]. The reference information in Table I suggests that most of the recent methods are in deep learning.
In this work, we compare our method with the following deep learning methods for low-light image enhancement: deep Retinex [24], zero-DCE [44], hybridNet [58], deep recursive band network (DRBN) [61], retinex-inspired unrolling with architecture search (RUAS) [62], RetinexDIP [63], enhancement-fusion iterative network (EFINet) [64], semantically contrastive learning for low-light image enhancement (SCL-LLE) [65], self-calibrated illumination (SCI) [66], semantic-guided zero-shot learning (SGZ) [67], unsupervised enhancement with generative adversarial network (UEGAN) [68], discrepant untrained network priors (DUNP) [43], generative diffusion prior (GDP) [42], NeurBR [41]. It is worth noting that we have included state-of-the-art artificial intelligence (AI) novel directions: generative adversarial network (GAN), diffusion method, zero-shot learning, etc.
We also compared two recent methods [37, 48] built on the idea of Poisson noise in low-light images. The authors in [48] introduced a regularizer based on the gamma-correction function and formulated it into a screened Poisson equation for the enhancement of low-light images. The iterative optimization algorithm in [37] employed regularization terms for the piecewise smooth prior of the illumination component and the Poisson noise distribution prior, which contributes to noise suppression in different noise intensity.
V-3 Visual Comparison
We show the visual results of our method along with the baseline in two different scenes with varying illumination levels. In our experiments, we have explicitly considered input images with four different levels of illumination (severe low-lightness as we move down from the first row to the fourth row). In Figure 3, we present the visual results of the enhancement of an AC image that is poorly illuminated at four different levels of low-lightness. Our method produces the best restored visual quality images without color distortion as shown in Figure 3(e). At first glance, we notice the evidence of color distortion by LIME and other methods shown in Figure 3. Moreover, the nature of color distortion is found to have a relation with the amount of low-lightness in the input image. Note that the color distortion gets worse with a very low level of illumination (as is the case for the last row), which is the case of extreme low-light conditions. Our method significantly outperforms the baseline methods in terms of producing color distortion-free images with high measure of naturalness.
For a deeper understanding of the color distortion using various methods, we studied the histogram relationship of {R, G, B} channels. An image with good color consistency is expected to exhibit very similar histograms for RGB colors. In Figure 4, we show the {R, G, B} histograms of the restored images using LIME, deep-retinex, zero-DCE, and our method PIRL (for low-light level 3 in Figure 3). Notice that our method gives the best color consistency resulting, no color distortion unlike other methods.
In Figure 5, we show the visual results of another low-light image suffered by four different levels of low-lightness. It is visually evident that the color distortion gets worse with a very low level of illumination, which is the case for extreme low-light conditions. In both Figures 3 and 5, we have used the same scene as input images but with varying exposure (along the rows). To be precise, as we move down from the first row to the fourth row, the exposure in the input image decreases. A lesser exposure image, for example, in the fourth row, has huge color shift towards green channel of the image. This prominent color shift in state-of-the-art methods increases as the exposure decreases in the image. For the same level of exposure, there is no color shift or distortion in the enhanced image produced by our method.
Based on exhaustive experiments, we conclude that as more an image is captured in lower exposure, more it will suffer from color shift or distortion when illumination enhancement is done while considering denoising along with it as separate task. When treating enhancement task as denoising only and considering an underexposure image as Poisson noise, it leads to attaining image enhancement without having a prominent color shift.
V-4 Quantitative Comparison
We evaluated the performance of various methods using three image quality metrics: peak signal-to-noise ratio (PSNR) [70], structural similarity index measure (SSIM) [71], and no-reference image quality evaluator (NIQE) [70]. With the availability of ground-truth in the LOL dataset, we computed the PSNR and SSIM values between the ground-truth and the output images (of each method). For both PSNR and SSIM, a higher value indicates better quality. For the NIQE metric, a lower value indicates better quality of the enhanced image. In Table I, we summarize the performance of low-light enhancement using our method and the baselines. Notice that our method PIRL achieves the second highest PSNR, (marginally) the highest SSIM, and the second best (lowest) NIQE values. We argue that overall our method achieves competitive performance with the baseline methods in terms of both quantitative measure and visual quality of the restored images. Experimental results confirm that our method is able to perform efficient low-light enhancement even under extreme low light conditions.
VI Conclusion
@@
In this work, we introduced a new deep learning method to address the challenging task of extreme low-light image enhancement and denoising. We demonstrated with experimental finding that many baseline methods suffer from color distortion when the illumination level of the scene/object is extremely poor. We introduced a novel image decomposition framework in contrast to Retinex model of image decomposition. Furthermore, our encoder-decoder-based deep learning method could produce superior recovery of extreme low-light images without any visible color distortion. Given that Poisson noise dominates in extreme low-light images because of limited photon counts, specialized Poisson denoising techniques are essential for preserving image quality and structural details.
In future research, we will focus on developing denoising models that can robustly handle mixed noise distributions while enabling computationally efficient (light-weight network) solutions for real-time feasibility. We also plan to integrate a Poisson denoising scheme directly on raw low-light images, which could offer a more lightweight and computationally efficient alternative. Leveraging raw image enhancement techniques, we would be able to improve performance utilizing unprocessed sensor data, enabling more accurate noise modeling. This direction of research will offer better structural preservation and enhanced visual quality in detail, while maintaining real-time applicability.